OpenStudy (anonymous):

f(x) = 3(x-4)^2 (x+4) what is the left and right hand behavior of this equation? What is the y intercept?

OpenStudy (anonymous):

y intercept = 3* 4^2*4 = 192 (192, 0) Both left and right raising up

OpenStudy (anonymous):

Do you know what the multiplicities of zero would be?

OpenStudy (phi):

(x-4)^2 will be positive for all real x (x+4) will be negative for x<-4. so for large negative x, the entire expression will become a large negative number. Going left, it tends to negative infinity. (down) Going right, it tends to positive infinity (up)